### v1

# Fine-grained Abnormality Prompt Learning for Zero-shot Anomaly Detection

Abstract
Current zero-shot anomaly detection (ZSAD) methods show remarkable success in prompting large pre-trained vision-language models to detect anomalies in a target dataset without using any dataset-specific training or demonstration. However, these methods are often focused on crafting/learning prompts that capture only coarse-grained semantics of abnormality, e.g., high-level semantics like ‘damaged’, ‘imperfect’, or ‘defective’ on carpet. They therefore have limited capability in recognizing diverse abnormality details with distinctive visual appearance, e.g., specific defect types like color stains, cuts, holes, and threads on carpet. To address this limitation, we propose FAPrompt, a novel framework designed to learn Fine-grained Abnormality Prompts for more accurate ZSAD. To this end, we introduce a novel compound abnormality prompting module in FAPrompt to learn a set of complementary, decomposed abnormality prompts, where each abnormality prompt is formed by a compound of shared normal tokens and a few learnable abnormal tokens. On the other hand, the fine-grained abnormality patterns can be very different from one dataset to another. To enhance their cross-dataset generalization, we further introduce a data-dependent abnormality prior module that learns to derive abnormality features from each query/test image as a sample-wise abnormality prior to ground the abnormality prompts in a given target dataset. Comprehensive experiments conducted across 19 real-world datasets, covering both industrial defects and medical anomalies, demonstrate that FAPrompt substantially outperforms state-of-the-art methods by at least 3%-5% AUC/AP in both image- and pixel-level ZSAD tasks. Code is available at https://github.com/mala-lab/FAPrompt.
1 Introduction
Anomaly Detection (AD) is a critical task in computer vision, aiming to identify instances that deviate significantly from the majority of data. It has a wide range of real-world applications, e.g., industrial inspection and medical imaging analysis (Pang et al., 2021; Cao et al., 2024). Traditional AD methods focus on learning specialized detectors with large training samples. Consequently, these methods often rely on application-specific, carefully curated datasets to train a detection model, making them inapplicable for application scenarios where such data access is not possible due to data privacy issue, or where the test data significantly differs from the training set due to substantial distribution shifts arising from new deployment environments or other natural variations in datasets. Zero-shot AD (ZSAD), which aims at learning generalist models for detecting anomalies in a target dataset without using any dataset-specific training or demonstration, has been recently emerging as a promising approach to address this limitation of traditional AD approaches.
In recent years, large pre-trained vision-language models (VLMs) such as CLIP (Radford et al., 2021) have demonstrated impressive zero/few-shot recognition capabilities across a broad range of vision tasks, including the ZSAD task (Chen et al., 2023b; Jeong et al., 2023; Deng et al., 2023; Zhou et al., 2024). To leverage VLMs for AD, the methods craft/learn text prompts to extract the textual semantic of normal/abnormal from VLMs for matching visual anomalies. These methods, such as WinCLIP (Jeong et al., 2023) and AnoVL (Deng et al., 2023), attempt to capture a range of abnormality semantics for better ZSAD by including a wide variety of pre-defined state-aware tokens (e.g., using ‘damaged’, ‘imperfect’, or ‘defective’ to depict defects on different objects like carpet) or domain-aware tokens (e.g., ‘industrial’, ‘manufacturing’, or ‘surface’) into the text prompts. Others (Zhou et al., 2024; 2022a; 2022b) employ learnable text prompts to extract more general-purpose features for representing the normal/abnormal class, such as AnomalyCLIP (Zhou et al., 2024). However, these approaches are focused primarily on coarse-grained semantics of abnormality and have limited capability in recognizing diverse abnormality details with distinctive visual appearance, e.g., specific defect types like color stains, cuts, holes, and threads on carpet, as shown in Fig. 1 Left, leading to failures in detecting anomalies that are different from the coarse-grained abnormal patterns. A recent approach AnomalyGPT (Gu et al., 2023) deals with this issue by using detailed text description of abnormal objects through an additional Large Language Model (LLM), but it requires the reference samples from the target data, which is a different task from ZSAD. It also heavily relies on costly human annotations for the detailed textual descriptions.
Figure 1: Left: FAPrompt vs. two related methods. Right: Their image-level ZSAD results in AUROC.
To tackle these issues, we propose a novel framework, namely FAPrompt, designed to learn Fine-grained Abnormality Prompts for more accurate ZSAD. In contrast to previous prompting methods, FAPrompt focuses on learning the prompts that can model diverse fine-grained abnormality semantics without requiring detailed human annotations or text descriptions, as illustrated in the bottom of Fig. 1 Left. To this end, in FAPrompt we introduce a novel Compound Abnormality Prompting module, namely CAP, to learn a set of complementary, decomposed abnormality prompts on top of a normal prompt, where each abnormality prompt is formed by a compound of the same tokens in the normal prompt and a few learnable abnormal tokens. The insight of this design is rooted from our observation that each abnormal pattern can be considered as some unexpected patterns overlaying on top of common normal patterns, e.g., color stains on normal texture of carpet. Such a compound prompting strategy enables the learning of different abnormality semantics easily while maintaining abnormality prompts in good proximity to the normal prompt. This helps avoid learning trivial abnormality prompts that are too far away from the normal prompt, lacking discriminability for distinguishing normal and abnormal samples.
On the other hand, the fine-grained abnormality patterns can be very different from one dataset to another. Thus, to achieve better cross-dataset generalization, the learned fine-grained abnormality prompts should be adaptive to any target testing datasets. We therefore further introduce a Data-dependent Abnormality Prior module, namely DAP, to enhance the cross-dataset generalizability of the abnormal tokens in CAP. It learns to derive abnormality features from each query/test image as a sample-wise abnormality prior to dynamically adapt the abnormality prompts in CAP to a given target dataset. This interaction between CAP and DAP enables the learning of abnormality prompts that have fine-grained semantics and are adaptive to different testing datasets, enabling better ZSAD across a wide range of image AD datasets, as shown in Fig. 1 Right.
Accordingly, we make the following main contributions.
  • We propose a novel ZSAD framework FAPrompt. Unlike existing methods that capture coarse-grained semantics of abnormality only, FAPrompt offers an effective approach for learning adaptive fine-grained abnormality semantics without any reliance on detailed human annotations or text descriptions.
  • To achieve this, we first introduce a novel Compound Abnormality Prompting module (CAP) in FAPrompt. It learns a small set of complementary, decomposed abnormality prompts on top of the normal prompt via a compounding normal-abnormal token design and an orthogonal constraint among the abnormality prompts.
  • We further introduce a Data-dependent Abnormality Prior module (DAP). It learns to select the most relevant abnormal features from anomaly images while refraining from normal images for adapting the fine-grained abnormalities learned in CAP to a given target dataset.
  • Comprehensive experiments on 19 diverse real-world industrial and medical image AD datasets show that FAPrompt significantly outperforms state-of-the-art ZSAD models by at least 3%-5% AUC/AP in both image- and pixel-level detection tasks.
2 Related Work
2.1 Conventional Anomaly Detection
There have been different types of AD approaches introduced over the years. In particular, one-class classification methods (Tax & Duin, 2004; Yi & Yoon, 2020; Bergman & Hoshen, 2020; Chen et al., 2022; Ruff et al., 2020) aim to compactly describe normal data using support vectors. Reconstruction-based methods (Akcay et al., 2019; Schlegl et al., 2019; Zavrtanik et al., 2021b; Yan et al., 2021; Zaheer et al., 2020; Zavrtanik et al., 2021a; Park et al., 2020; Hou et al., 2021; Xiang et al., 2023; Liu et al., 2023; Yao et al., 2023b; a) train models to reconstruct normal images, with anomalies identified through higher reconstruction errors. Distance-based methods (Pang et al., 2018; Defard et al., 2021; Cohen & Hoshen, 2020; Roth et al., 2022) detect anomalies by measuring the distance between the test image and normal images. Knowledge distillation methods (Deng & Li, 2022; Bergmann et al., 2020; Salehi et al., 2021; Wang et al., 2021; Cao et al., 2023; Tien et al., 2023; Zhang et al., 2023) focus on distilling normal patterns from pre-trained models and detecting anomalies by comparing discrepancies between the distilled and original features. However, these methods often rely on application-specific datasets to train the detection model, limiting their applicability in real-world scenarios where data access is restricted due to privacy concerns, proprietary restrictions, or resource constraints. Also, these approaches tend to struggle when there is a significant difference between the distribution of the training and test data.
2.2 Zero-Shot Anomaly Detection
ZSAD has been made possible due to the development of large pre-trained foundation models, such as vision-language models (VLMs). CLIP (Radford et al., 2021) has been widely used as a VLM to enable ZSAD on visual data (Jeong et al., 2023; Zhou et al., 2024; Deng et al., 2023; Chen et al., 2023a). CLIP-AC adapts CLIP for ZSAD by using text prompts designed for the ImageNet dataset as in  (Radford et al., 2021). By using manually defined textual prompts specifically designed for industrial AD dataset, WinCLIP (Jeong et al., 2023) achieves better ZSAD performance compared to CLIP-AC, but it often does not generalize well to non-defect AD datasets. APRIL-GAN (Chen et al., 2023a) adapts CLIP to ZSAD through tuning some additional linear layers with annotated auxiliary AD data. AnoVL (Deng et al., 2023) introduces domain-aware textual prompts and test time adaptation in CLIP to enhance the ZSAD performance. AnomalyCLIP (Zhou et al., 2024) employs learnable, object-agnostic textual prompts to extract more general-purpose text features for the normal and abnormal classes. All these methods are focused on crafting/learning prompts that capture only coarse-grained semantics of abnormality, failing to detect anomalies that exhibit different patterns from these coarse abnormal patterns. There are a number of other studies leveraging CLIP for AD, but they are designed for empowering few-shot (Gu et al., 2023; Zhu & Pang, 2024) or conventional AD task (Joo et al., 2023; Wu et al., 2024a; b; c).
Figure 2: Overview of FAPrompt. It consists of two novel modules, including the Compound Abnormality Prompting (CAP) module and the Data-dependent Abnormality Prior (DAP) module detailed in the top-right and bottom-right corners respectively. CAP is devised to learn fine-grained abnormality semantics without relying on detailed human annotations or text descriptions, while DAP is designed to adaptively select the most abnormal features from each query/test image as a sample-wise abnormality prior to enhance the cross-dataset generalizability of the abnormality prompts in CAP.
3 Methodology
3.1 Preliminaries
Problem Statement. Let $\mathcal{D}_{train}=\{X_{train},Y_{train}\}$ denote an auxiliary training dataset consisting of both normal and anomalous samples, where $X_{train}=\{x_{i}\}_{i=1}^{N}$ is a set of $N$ images and $Y_{train}=\{y_{i},\mathbf{G}_{i}\}_{i=1}^{N}$ contains the corresponding ground truth labels and pixel-level anomaly masks. Each image $x_{i}$ is labeled by $y_{i}$, where $y_{i}=0$ indicates a normal image and $y_{i}=1$ signifies an anomalous one. The anomaly mask $\mathbf{G}_{i}$ provides pixel-level annotation of $x_{i}$. During the testing phase, we are presented with a collection of target datasets, $\mathcal{T}=\{\mathcal{D}_{test}^{1},\mathcal{D}_{test}^{2},\cdots,\mathcal{D}_{test}^{t}\}$, where each $\mathcal{D}_{test}^{j}=\{X_{test}^{j},Y_{test}^{j}\}$ is a test set from a target application dataset that have different normal and abnormal samples from those in the training data $\mathcal{D}_{train}$. The goal of ZSAD is to develop models on the auxiliary dataset $\mathcal{D}_{train}$, with the ability to generalize to detect anomalies in different test sets in $\mathcal{T}$. Particularly, given an input RGB image $x\in\mathbb{R}^{h\times w\times 3}$ from $\mathcal{D}_{train}$, with $h$ and $w$ respectively representing the height and width of $x$, a ZSAD model is required to output both an image-level anomaly score $s_{x}\in\mathbb{R}$ and a pixel-level anomaly map $\mathbf{\mathcal{M}}_{x}\in\mathbb{R}^{h\times w}$. The image-level anomaly score $s_{x}$ provides a global assessment of whether the image is anomalous, while the pixel-level anomaly map $\mathbf{\mathcal{M}}_{x}$ indicates the likelihood of each pixel being anomalous. Both $s_{x}$ and the values in $\mathbf{\mathcal{M}}_{x}$ lie in [0, 1], where a larger value indicates a higher probability of being abnormal.
VLM Backbone. To enable accurate ZSAD, large pre-trained VLMs are typically required. Following existing approaches (Chen et al., 2023b; Deng et al., 2023; Jeong et al., 2023; Zhou et al., 2024), the pre-trained CLIP (Radford et al., 2021) is used in our study, which comprises a visual encoder $f_{v}(\cdot)$ and a text encoder $f_{t}(\cdot)$, where the visual and text representations are well-aligned through pre-training on web-scaled text-image pairs.
3.2 Overview of FAPrompt
In this work, we propose a ZSAD framework FAPrompt to learn adaptive fine-grained abnormality semantics without any reliance on detailed human annotations or text descriptions. Fig 2 illustrates the overall framework of FAPrompt that consists of two novel modules, including Compound Abnormality Prompting module (CAP) and Data-dependent Abnormality Prior module (DAP). To be more specific, the proposed CAP module is devised to specify the design of fine-grained abnormality prompts. The key characteristic of CAP is to obtain the abnormality prompts via a compound prompting method, where we have one normal prompt and multiple abnormality prompts are added on top of it. These normal and abnormal text prompts are then processed by the CLIP’s text encoder $f_{t}(\cdot)$ to generate the corresponding normal and abnormal text embeddings, respectively. For a given image $x$, FAPrompt extracts both an image token embedding $f_{v}(x)$ and a set of patch token embeddings $\mathbf{F}_{v}\in\mathbb{R}^{l\times d}$, with $l$ and $d$ respectively representing the length and dimensionality of $\mathbf{F}_{v}$. The prompts are then learned using $\mathcal{D}_{train}$ based on the similarity between the image and text embeddings, where the fine-grained abnormality prompts are aggregated into an abnormality prompt prototype before its use in similarity calculation. Further, the DAP module is introduced to improve cross-dataset generalization capability of the fine-grained abnormality prompts. DAP derives the most relevant abnormality features based on the given query/test image $x$, serving as a sample-wise abnormality prior to dynamically adapt the abnormality prompts in CAP to the characteristics of a given target dataset. During training, the original parameters of CLIP remain frozen, and only the attached learnable tokens in the text encoder layers, along with the normal and fine-grained abnormality prompts, are optimized. Below we present these modules in detail.
3.3 Compound Abnormality Prompt Learning
Learning Fine-grained Abnormalities via Compound Normal and Abnormal Tokens. Previous approaches that rely on coarse-grained learnable text prompts fail to capture the fine-grained abnormality semantics for detecting diverse anomalies across different datasets. To address this, we propose the novel CAP module. The core insight is that abnormal samples typically exhibit different magnitude of deviation from their normal counterparts while still belonging to the same class. CAP models this by learning a set of complementary, decomposed abnormality prompts built on a shared normal prompt. Following previous work (Zhou et al., 2024), a set of learnable normal tokens and the fixed token ‘object’ are concatenated to define the normal text prompt $\mathcal{P}^{n}$. For the abnormality prompt, CAP aims to learn a small set of prompts of complementary semantics, denoted as $\mathcal{P}^{a}=\{\mathcal{P}^{a_{1}},\mathcal{P}^{a_{2}},...\mathcal{P}^{a_{K}}\}$, where each $\mathcal{P}^{a_{i}}$ is formed by a compound of the same tokens in the normal prompt $\mathcal{P}^{n}$ and a few learnable abnormal tokens. Formally, the normal and abnormality prompts can be defined as:
$\mathcal{P}^{n}=[V_{1}][V_{2}]...[V_{E}][object],$ (1)
$\mathcal{P}^{a}=\{\mathcal{P}^{a_{1}},\mathcal{P}^{a_{2}},...\mathcal{P}^{a_{K}}\},$
$with\ \mathcal{P}^{a_{i}}$ $=[V_{1}][V_{2}]...[V_{E}][A_{1}^{i}][A_{2}^{i}]...[A_{E^{\prime}}^{i}][object],$
where $\{V_{1},V_{2},...V_{E}\}$ and $\{A_{1}^{i},A_{2}^{i},...A_{E^{\prime}}^{i}\}_{i=1}^{K}$ are learnable normal and abnormal tokens, respectively. This compound prompting strategy enables the learning of different abnormality semantics easily while maintaining abnormality prompts in good proximity to the normal prompt, supporting the learning of non-trivial, semantically-meaningful abnormality prompts.
Learning Complementary Abnormality Prompts. To capture complementary fine-grained abnormalities and reduce redundant information captured by the abnormality prompts, it is essential to maximize the diversity among the fine-grained abnormalities. A straightforward approach would be to train distinct abnormal prompts on separate, annotated subsets with samples from different anomalous types. However, this would require extensive human annotations. To address this issue, we propose to add an orthogonal constraint loss $\mathcal{L}_{oc}$ into the abnormality prompts in CAP as a alternative method to encourage this diversity. Formally, the objective for this can be formulated as:
$\mathcal{L}_{oc}=\sum_{i,j\in K;i\neq j}abs(\frac{f_{t}(\mathcal{P}^{a_{i}})\cdot f_{t}(\mathcal{P}^{a_{j}})}{||f_{t}(\mathcal{P}^{a_{i}})||\times||f_{t}(\mathcal{P}^{a_{j}})||}),$ (2)
where the text encoder $f_{t}(\cdot)$ is used to extract the embeddings of the abnormality prompts, []delimited-[][$\cdot$][ ⋅ ] denotes inner product, $abs(\cdot)$ returns the absolute value, and |||||$\cdot$||| | ⋅ | | indicates the norm of vectors.
To provide more representative embedding for the fine-grained abnormalities, we compute the prototype of the multiple abnormality prompt embeddings as the final fine-grained abnormality embedding via $\mathbf{F}_{a}=\frac{1}{|\mathcal{P}^{a}|}\sum_{\mathcal{P}^{a_{i}}\in\mathcal{P}^{a}}f_{t}(\mathcal{P}^{a_{i}})$. The normal text prompt embedding is $\mathbf{F}_{n}=f_{t}(\mathcal{P}^{n})$.
3.4 Learning to Select Data-dependent Abnormality Prior
One issue in ZSAD is that the fine-grained abnormality patterns can be very different from the auxiliary dataset to test datasets. In addition to the learning of a set of complementary fine-grained abnormality prompts, it is important to ensure that the learned fine-grained abnormality patterns are generalized to target testing datasets. Inspired by the instance-conditional information design in CoCoOp (Zhou et al., 2022a), we introduce the DAP module to enhance the cross-dataset generalizability of the abnormal tokens in CAP by adaptively selecting the embeddings of the most abnormal regions to serve as a sample-wise abnormality prior for each image input. Particularly, given a query/test image $x$, DAP selects the most abnormal image patches as the abnormality prior to be fed into CAP for assisting the abnormality prompt learning. It achieves this by picking the top $M$ patches whose token embeddings are most similar to the abnormality prompt prototype $\mathbf{F}_{a}$:
$\mathbf{S}_{x}^{a}(i,j)=\frac{\exp(\mathbf{F}_{v}(i,j)\mathbf{F}_{a}^{\intercal})}{\exp(\mathbf{F}_{v}(i,j)\mathbf{F}_{n}^{\intercal})+\exp(\mathbf{F}_{v}(i,j)\mathbf{F}_{a}^{\intercal})},$ (3)
where $[\cdot]^{\intercal}$ denotes a transpose operation, $\mathbf{F}_{v}(i,j)$ is the token embedding of the patch centered at $(i,j)$ and $\mathbf{S}_{x}^{a}(i,j)$ is a patch-level anomaly score. The corresponding normal scores can be calculated via $\mathbf{S}_{x}^{n}(i,j)$ using the similarity to $\mathbf{F}_{n}$ in the numerator in Eq. 3.
Let $\textbf{p}_{x}=\{p_{1},p_{2},...p_{M}\}$ be the top $M$ patch embeddings of $x$, FAPrompt then adds additional learnable layers $\psi(\cdot)$, namely abnormality prior network, to model the sample-wise abnormality prior based on $\textbf{p}_{x}$. This prior $\Omega_{x}=\psi(\textbf{p}_{x})$ is then incorporated as data-dependent abnormal features into the learnable abnormal tokens of the abnormality prompts in CAP to dynamically adapt the learned fine-grained abnormalities to a given target dataset, with each individual abnormality prompt refined as follows:
$\mathcal{\hat{P}}^{a_{i}}=[V_{1}][V_{2}]...[V_{E}][A_{1}^{i}\oplus\Omega_{x}][A_{2}^{i}\oplus\Omega_{x}]...[A_{E^{\prime}}^{i}\oplus\Omega_{x}][object],$ (4)
where $\Omega_{x}$ is a vector-based prior of the same dimensionality as the abnormal tokens and $\oplus$ denotes element-wise addition. Thus, the abnormality prompt set is updated as $\mathcal{\hat{P}}^{a}=\{\mathcal{\hat{P}}^{a_{1}},\mathcal{\hat{P}}^{a_{2}},...\mathcal{\hat{P}}^{a_{K}}\}$, and the abnormality prompt prototype can be accordingly refined as $\mathbf{\hat{F}}_{a}=\frac{1}{|\mathcal{\hat{P}}^{a}|}\sum_{\mathcal{\hat{P}}^{a_{i}}\in\mathcal{\hat{P}}^{a}}f_{t}(\mathcal{\hat{P}}^{a_{i}})$.
The goal of DAP is to introduce sample-wise abnormality information. However, there is no abnormality from the top $M$ patches of normal images, and thus, simply applying the prior $\Omega_{x}$ to normal images would introduce noise into the learnable abnormal tokens, damaging the learning of fine-grained abnormalities. To address this issue, we propose an abnormality prior learning loss $\mathcal{L}_{prior}$ to enforce that $\Omega_{x}$ is the features mapped from the most abnormal $M$ patches if $x$ is an abnormal image, while it is minimized to be a null vector if it is a normal image. Formally, $\mathcal{L}_{prior}$ can be defined as follows:
$\mathcal{L}_{prior}=\sum_{y_{x}=0}\sum_{\omega\in\Omega_{x}}\omega_{x}^{2},$ (5)
where $\omega$ is an entry of $\Omega_{x}$.
3.5 Training and Inference
[…]
where I is a full-one matrix, $\mathcal{L}_{Focal}(\cdot)$ and $\mathcal{L}_{Dice}(\cdot)$ denote a focal loss (Lin et al., 2017) and a Dice loss (Li et al., 2019b), respectively. To ensure the accuracy of locating the top abnormal features in DAP, we apply the same learning objective to optimize the segmentation maps $\mathbf{\mathcal{M}^{n}}\in\mathbb{R}^{h\times w}$ and $\mathbf{\mathcal{M}^{a}}\in\mathbb{R}^{h\times w}$, which are derived from the normality-oriented scores ${S_{x}^{n}}$ and abnormality-oriented scores ${S_{x}^{a}}$, respectively.
For image-level supervision, FAPrompt first computes the probability of the query image $x$ being classified as abnormal based on its cosine similarity to the two prompt embeddings $\mathbf{\hat{F}}_{a}$ and $\mathbf{F}_{n}$:
$s_{a}(x)=\frac{\exp(f_{v}(x)\mathbf{\hat{F}}_{a}^{\intercal})}{\exp(f_{v}(x)\mathbf{F}_{n}^{\intercal})+\exp(f_{v}(x)\mathbf{\hat{F}}_{a}^{\intercal})}.$ (8)
The final image-level anomaly score is then defined as the average of this image-level score and the maximum pixel-level anomaly score derived from the anomaly score maps:
$s(x)=\frac{1}{2}(s_{a}(x)+s_{a}^{\prime}(x)),$ (9)
where $s_{a}^{\prime}(x)=\frac{1}{2}\left(\max({S_{x}^{a}})+\max({\hat{S}}_{x}^{a})\right)$ represents the average of the maximum anomaly scores from ${S_{x}^{a}}$ and ${\hat{S}_{x}^{a}}$. Following previous methods (Zhu & Pang, 2024; Chen et al., 2023a; Zhou et al., 2024; Jeong et al., 2023), $s_{a}^{\prime}(x)$ is treated as a complementary anomaly score to $s_{a}(x)$ and incorporated into Eq. 9, as $s_{a}^{\prime}(x)$ are helpful for detecting local abnormal regions. The image-level anomaly score $s(x)$ is then optimized by minimizing the following loss on $X_{train}$:
$\mathcal{L}_{global}=\frac{1}{N}\sum_{x\in X_{train}}\mathcal{L}_{b}(s(x),y_{x}),$ (10)
where $\mathcal{L}_{b}$ is specified by a focal loss function due to the class imbalance in $X_{train}$. Overall, FAPrompt is optimized by minimizing the following combined loss, which integrates both local and global objectives, along with the two constraints from the CAP and DAP modules:
$\mathcal{L}=\mathcal{L}_{local}+\mathcal{L}_{global}+\mathcal{L}_{prior}+\mathcal{L}_{oc},$ (11)
[…]
where $\ominus$ is element-wise subtraction. The image-level anomaly score $s_{x^{\prime}}$ is computed using Eq. 9.
4 Experiments
Datasets. To verify the effectiveness of FAPrompt, we conduct extensive experiments across 19 publicly available datasets, including nine popular industrial defect inspection datasets on varying products/objects (MVTecAD (Bergmann et al., 2019), VisA (Zou et al., 2022), DAGM (Wieler & Hahn, 2007), DTD-Synthetic (Aota et al., 2023), AITEX (Silvestre-Blanes et al., 2019), SDD (Tabernik et al., 2020), BTAD (Mishra et al., 2021), MPDD (Jezek et al., 2021), and ELPV(Deitsch et al., 2019)) and ten medical anomaly detection datasets on different organs like brain, fundus, colon, skin and thyroid (BrainMRI (Salehi et al., 2021), HeadCT (Salehi et al., 2021), LAG (Li et al., 2019a), Br35H (Hamada, ), CVC-ColonDB (Tajbakhsh et al., 2015), CVC-ClinicDB (Bernal et al., 2015), Kvasir (Jha et al., 2020), Endo (Hicks et al., 2021), ISIC (Gutman et al., 2016), TN3K (Gong et al., 2021)) (see Appendix A for details about the datasets).
Table 1: Image-level ZSAD results (AUROC, AP) on 13 AD datasets. The best and second-best results are respectively highlighted in red and blue. The results for MVTecAD, VisA, DAGM, DTD-Synthetic, BTAD, and MPDD are averaged performance across their multiple data subsets (see Appendix D for breakdown results).
Data Type Dataset Handcrafted Text Prompts Learnable Text Prompts
CLIP CLIP-AC WinCLIP APRIL-GAN AnoVL CoOp CoCoOp AnomalyCLIP FAPrompt
Industrial MVTecAD (74.1, 87.6) (71.5, 86.4) (91.8, 96.5) (86.2, 93.5) (92.5, 96.7) (88.8, 94.8) (71.8, 84.9) (91.5, 96.2) (91.9, 95.7)
VisA (66.4, 71.4) (65.0, 70.2) (78.8, 81.4) (78.0, 81.4) (79.2, 81.7) (62.8, 68.1) (78.1, 82.3) (82.1, 85.4) (84.5, 86.8)
SDD (95.5, 87.9) (94.7, 77.9) (94.0, 87.2) (97.5, 93.4) (95.3, 91.3) (96.8, 90.0) (89.9, 50.4) (98.1, 93.4) (98.6, 95.9)
BTAD (34.5, 52.5) (51.0, 62.1) (68.2, 70.9) (73.6, 68.6) (80.3, 73.1) (66.8, 77.4) (48.4, 53.9) (88.3, 87.3) (92.0, 92.2)
MPDD (54.3, 65.4) (56.2, 66.0) (63.6, 69.9) (73.0, 80.2) (68.9, 71.9) (55.1, 64.2) (61.0, 69.1) (77.0, 82.0) (80.6, 83.3)
AITEX (71.0, 45.7) (71.5, 46.7) (73.0, 54.7) (57.6, 41.3) (72.5, 55.4) (66.2, 39.0) (48.6, 37.8) (62.2, 40.4) (71.9, 53.2)
DAGM (79.6, 59.0) (82.5, 63.7) (91.8, 79.5) (94.4, 83.8) (89.7, 76.3) (87.5, 74.6) (96.3, 85.5) (97.5, 92.3) (98.9, 95.7)
DTD-Synthetic (71.6, 85.7) (66.8, 83.2) (93.2, 92.6) (86.4, 95.0) (94.9, 97.3) (83.1, 91.9) (84.1, 92.9) (93.5, 97.0) (95.9, 98.3)
ELPV (59.2, 71.7) (69.4, 80.2) (74.0, 86.0) (65.5, 79.3) (70.6, 83.0) (73.0, 86.5) (78.4, 89.2) (81.5, 91.3) (83.5, 92.0)
Medical BrainMRI (73.9, 81.7) (80.6, 86.4) (86.6, 91.5) (89.3, 90.9) (88.7, 91.3) (61.3, 44.9) (78.2, 86.7) (90.3, 92.2) (95.5, 95.6)
HeadCT (56.5, 58.4) (60.0, 60.7) (81.8, 80.2) (89.1, 89.4) (81.6, 84.2) (78.4, 78.8) (80.3, 73.4) (93.4, 91.6) (94.8, 93.5)
LAG (58.7, 76.5) (58.2, 76.9) (59.2, 74.8) (73.6, 84.8) (65.1, 78.0) (69.6, 82.9) (72.6, 84.7) (74.3, 84.9) (75.6, 85.4)
Br35H (78.4, 78.8) (82.7, 81.3) (80.5, 82.2) (93.1, 92.9) (88.4, 88.9) (86.0, 87.5) (85.7, 89.1) (94.6, 94.7) (97.8, 97.5)
To assess the ZSAD performance, the models are trained on the MVTecAD dataset by default and evaluated on the test sets of other datasets without any further training or fine-tuning. We obtain the ZSAD results on MVTecAD by changing the training data to the VisA dataset.
Competing Methods and Evaluation Metrics. We compare our method, FAPrompt, with several state-of-the-art (SotA) methods, including five handcrafted text prompt-based methods – raw CLIP (Radford et al., 2021), CLIP-AC and WinCLIP (Jeong et al., 2023), APRIL-GAN (Chen et al., 2023a), and AnoVL (Deng et al., 2023) – and three learnable text prompt-based methods – CoOp (Zhou et al., 2022b), CoCoOp (Zhou et al., 2022a), and AnomalyCLIP (Zhou et al., 2024). As for evaluation metrics, we follow previous works (Jeong et al., 2023; Zhou et al., 2024) and use two popular metrics: AUROC (Area Under the Receiver Operating Characteristic) and average precision (AP) to assess the image-level AD performance; for pixel-level AD performance, we employ AUROC and Area under per region overlap (PRO) to provide a more detailed analysis.
Implementation Details. Following previous approaches (Zhou et al., 2024; Chen et al., 2023a), we implement FAPrompt using the same CLIP implementation, OpenCLIP (Ilharco et al., 2021), using the publicly available pre-trained VIT-L/14@336px backbone. The parameters of both the visual and text encoders in CLIP are kept frozen. By default, learnable token embeddings are attached to the first nine layers of the text encoder, with a token length of four for each layer. The lengths of the learnable normal and abnormal text prompts are respectively set to five and two by default. The number of fine-grained abnormality prompts and selected patch tokens in the DAP module are both set to ten. We use the Adam optimizer with an initial learning rate of 1e-3. Further implementation details for FAPrompt and the competing methods are provided in Appendix B.
4.1 Main Results
Image-level ZSAD Performance. Table 1 presents the image-level ZSAD results of FAPrompt, compared to eight SotA methods across 13 AD datasets, including nine industrial defect AD datasets and four medical AD datasets. The results show that FAPrompt significantly outperforms the SotA models across almost all datasets. On average, compared to the best competing methods, it achieves up to 3.7% AUROC and 4.9% AP on industrial AD datasets and 5.2% AUROC and 3.4% AP on medical AD datasets. In particular, the weak performance of CLIP and CLIP-AC can be attributed to its over-simplified text prompt design. By utilizing more carefully designed handcrafted prompts, WinCLIP achieves better results than CLIP and CLIP-AC while preserving the training-free nature. APRIL-GAN and AnoVL improve over WinCLIP by using additional learnable layers and/or domain-aware tokens within the textual prompts. However, they heavily rely on sensitive handcrafted textual prompts and capture mainly coarse-grained semantics of abnormality, leading to poor performance when faced with anomalies that does not fit well to the pre-defined text descriptions, e.g., BTAD, MPDD, BrainMRI, HeadCT, and Br35H.
As for text prompt learning methods, CoOp and CoCoOp are designed for general vision tasks, i.e., discriminating different objects, so they have weak capability in capturing the differences between normality and abnormality on the same object. AnomalyCLIP significantly improves performance by learning object-agnostic textual prompts for AD, demonstrating strong generalization capabilities across diverse datasets. However, AnomalyCLIP overlooks fine-grained abnormality details. FAPrompt overcomes this limitation via its two novel modules, CAP and DAP.
Table 2: Pixel-level ZSAD results (AUROC, PRO) on 14 AD datasets. The best and second-best results are respectively highlighted in red and blue. Note that medical datasets in Table 1 do not have pixel-level ground truth. Thus, different medical datasets are used here. Detailed breakdown results for MVTecAD, VisA, DAGM, DTD-Synthetic, BTAD, and MPDD can be found in Appendix D.
Data Type Dataset Handcrafted Text Prompts Learnable Text Prompts
CLIP CLIP-AC WinCLIP APRIL-GAN AnoVL CoOp CoCoOp AnomalyCLIP FAPrompt
Industrial MVTecAD (38.4, 11.3) (38.2, 11.6) (85.1, 64.6) (87.6, 44.0) (89.8, 76.2) (33.3, 6.6) (86.7, 79.6) (91.1, 81.4) (90.6, 83.3)
VisA (46.6, 14.8) (47.8, 17.2) (79.6, 56.8) (94.2, 86.8) (89.9, 71.2) (24.1, 3.8) (93.6, 86.7) (95.5, 87.0) (95.9, 87.5)
SDD (28.4, 5.1) (33.5, 7.6) (95.9, 78.4) (93.0, 84.6) (97.9, 82.6) (91.8, 81.7) (93.7, 85.0) (98.1, 95.2) (98.3, 93.6)
BTAD (30.6, 4.4) (32.8, 8.3) (72.7, 27.3) (60.8, 25.0) (93.2, 62.8) (28.6, 3.8) (86.1, 72.0) (94.2, 74.8) (95.6, 75.2)
MPDD (62.1, 33.0) (58.7, 29.1) (76.4, 48.9) (94.1, 83.2) (84.0, 61.0) (15.4, 2.3) (95.2, 84.2) (96.5, 87.0) (96.5, 87.9)
AITEX (53.2, 15.3) (47.3, 11.8) (62.5, 41.5) (78.2, 68.8) (59.2, 49.1) (67.7, 54.9) (52.1, 56,9) (83.0, 66.5) (82.0, 62.6)
DAGM (28.2, 2.9) (32.7, 4.8) (87.6, 65.7) (82.4, 66.2) (92.0, 78.8) (17.5, 2.1) (82.8, 75.1) (95.6, 91.0) (98.3, 95.4)
DTD-Synthetic (33.9, 12.5) (23.7, 5.5) (83.9, 57.8) (95.3, 86.9) (97.5, 90.4) (55.8, 36.0) (93.7, 83.7) (97.9, 92.3) (98.3, 93.1)
Medical CVC-ColonDB (49.5, 15.8) (49.5, 11.5) (70.3, 32.5) (78.4, 64.6) (77.9, 49.8) (40.5, 2.6) (79.1, 69.7) (81.9, 71.3) (84.6, 74.7)
CVC-ClinicDB (47.5, 18.9) (48.5, 12.6) (51.2, 13.8) (80.5, 60.7) (82.1, 55.0) (34.8, 2.4) (83.4, 68.8) (82.9, 67.8) (84.7, 70.1)
Kvasir (44.6, 17.7) (45.0, 16.8) (69.7, 24.5) (75.0, 36.2) (72.5, 28.2) (44.1, 3.5) (79.1, 38.6) (78.9, 45.6) (81.2, 47.8)
Endo (45.2, 15.9) (46.6, 12.6) (68.2, 28.3) (81.9, 54.9) (80.5, 47.7) (40.6, 3.9) (83.1, 59.0) (84.1, 63.6) (86.4, 67.2)
ISIC (33.1, 5.8) (36.0, 7.7) (83.3, 55.1) (89.4, 77.2) (90.6, 79.8) (51.7, 15.9) (81.9, 68.9) (89.7, 78.4) (90.9, 81.2)
TN3K (42.3, 7.3) (35.6, 5.2) (70.7, 39.8) (73.6, 37.8) (80.9, 50.5) (34.0, 9.5) (72.4, 41.0) (81.5, 50.4) (84.5, 54.1)
Pixel-level ZSAD Performance. We also compare the pixel-level ZSAD results of our FAPrompt with SotA methods across 14 AD datasets in Table 2. Similar observations can be derived as the image-level results. In particular, CLIP and CLIP-AC are the weakest among the handcrafted text prompt-based methods, primarily due to inappropriate text prompt designs. With better prompt engineering (and adaptation to AD in some cases), WinCLIP, APRIL-GAN, and AnoVL demonstrate better performance. For the learnable text prompt approaches, CoOp shows poor performance due to overfitting on the adaptation dataset, while CoCoOp mitigates this limitation by introducing instance-conditional information, achieving substantial improvement over CoOp and competitive performance to AnomalyCLIP. FAPrompt demonstrates superior performance in identifying a wide range of pixel-level anomalies, significantly outperforming SotA models across nearly all datasets. It surpasses the best competing methods by up to 2.7% AUROC and 4.4% AP on the industrial AD datasets, and by 3.0% AUROC and 3.7% AP on the medical AD datasets. This demonstrates the effectiveness of the fine-grained abnormality prompt in FAPrompt that adaptively capture detailed abnormality semantics in different datasets.
Performance of Learning Complementary Abnormalities. To assess the complementarity of the abnormality prompts learned by FAPrompt, we empirically evaluate the discriminability of each abnormality prompt and its difference to the rest of
Figure 3: t-SNE visualization of prompt-wise anomaly scores on BTAD (01).
other prompts. To this end, we calculate the patch-level anomaly scores $\mathbf{S}_{x}^{a}$ of an image based on the similarity of its patch token embeddings to each individual abnormal prompt embedding (rather than the prototype of the abnormal prompt embeddings) in CAP, and subsequently project the anomaly scores of each sample into a two-dimensional space via t-SNE. As depicted in Fig. 3 on an exemplar dataset, two key observations can be derived: i) despite having slight overlapping, the normal and abnormal samples are distributed into a different group for each individual abnormality prompt, indicating the learning of different abnormal patterns per abnormality prompt; and ii) there is clear separation between normal and abnormal samples for the use of each abnormality prompt in anomaly scoring, indicating the good discriminability of each prompt learned in FAPrompt. Similar patterns can be found in other datasets (see Appendix C.3).
To learn more abnormalities, one straightforward solution is to learn multiple abnormality prompts by using an ensemble of 10 abnormality prompts with each learned by AnomalyCLIP tuned on the auxiliary dataset with a different random seed. We argue that this simple strategy leads to the learning of highly redundant abnormality prompts, rather than the complementary prompts learned in FAPrompt. This is justified by our results of comparing FAPrompt to this ensemble method in Table 4.1. It is clear that the ensemble method can improve AnomalyCLIP to some extent,
Table 3: FAPrompt vs. AnomalyCLIP ensemble.
Model Industrial Datasets Medical Datasets
Image-level Pixel-level Image-level Pixel-level
AnomalyCLIP (85.0, 83.6) (94.4, 84.8) (87.7, 90.6) (83.2, 62.9)
AnomalyCLIP Ensemble (85.5, 84.0) (94.7, 85.0) (89.3, 91.3) (83.2, 62.4)
FAPrompt (88.2, 87.2) (95.0, 85.0) (90.9, 93.0) (85.4, 65.9)
but its abnormality prompts are much less effective than FAPrompt in both image- and pixel-level ZSAD results, despite that no ensemble strategy is used in FAPrompt. This showcases the effectiveness of the abnormality prompts learned in FAPrompt in capturing the fine-grained abnormality details.
4.2 Ablation Study
[…]
Table 4: Image-level (AUROC, AP) and pixel-level (AUROC, PRO) results of ablation study.
Model Industrial Datasets Medical Datasets
Image-level Pixel-level Image-level Pixel-level
Base (85.0, 83.6) (94.4, 84.8) (87.7, 90.6) (83.2, 62.9)
+ CAP (88.1, 87.0) (94.6, 83.9) (90.6, 93.1) (83.8, 63.8)
[…]
+ DAP (86.9, 85.2) (94.8, 84.9) (90.2, 92.3) (84.6, 64.8)
[…]
FAPrompt (88.2, 87.2) (95.0, 85.0) (90.9, 93.0) (85.4, 65.9)
[…]
Hyperparameter Sensitivity Analysis. We analyze the sensitivity of two key hyperparameters of FAPrompt in terms of image-level performance in Fig. 4, including the number of abnormality prompts $K$ in CAP and the number of selected patch tokens $M$ in DAP (similar results can be found for pixel-level results in Appendix C.4). In particular, the performance gets improved with increasing $K$, typically peaking at $K=10$. The performance
Figure 4: Averaged results on industrial and medical datasets with varying $K$ and $M$.
may slightly declines when $K$ is chosen beyond 10. This suggests that while increasing the number of prompts helps capture a wider range of abnormalities, too large $K$ values may introduce noise or redundancy into the prompts. As for the number of selected tokens, $M$, the performance exhibits a similar pattern, with the best performance obtained at a medium value. This suggests that selecting too many abnormal patch candidates may introduce noise or less relevant patches into CAP, leading to the learning of less effective fine-grained anomalities.
Additionally, we also evaluate the sensitivity of the length of learnable normal and abnormal tokens $\{E,E^{\prime}\}$ in CAP module. The Image-level (‘I-AUROC’ and ‘I-AP’)
Figure 5: Averaged results of FAPrompt with varying prompt sizes of $(E,E^{\prime})$.
and pixel-level (‘P-AUROC’ and ‘P-PRO’) ZSAD results are shown in Fig. 5. Overall, the setting of (5, 2) works best for both industrial and medical AD, yielding strong ZSAD performance. Longer prompt lengths, such as (10, 4), can introduce more complexity without clear performance improvement, particularly in pixel-level performance. Using shorter prompt lengths, e.g., the setting of (2, 1), lacks sufficient capacity to support the ZSAD task, leading to consistently weaker performance.
5 Conclusion
In this paper, we propose FAPrompt, a novel framework designed to enhance CLIP’s performance in ZSAD by learning adaptive fine-grained abnormality semantics. FAPrompt introduces a Compound Abnormality Prompting (CAP) module that generates complementary abnormality prompts without relying on exhausting human annotations. Additionally, it incorporates a Data-dependent Abnormality Prior (DAP) module, which refines these prompts to improve cross-dataset generalization. The interaction between CAP and DAP enables the model to learn adaptive fine-grained abnormality semantics. Extensive experiments on 19 datasets demonstrate that FAPrompt significantly outperforms state-of-the-art ZSAD methods.
Appendix A Dataset Details
A.1 Data Statistics of Training and Testing
We conduct extensive experiments on 19 real-world Anomaly Detection (AD) datasets, including nine industrial defect inspection datasets (MVTecAD (Bergmann et al., 2019), VisA (Zou et al., 2022), DAGM (Wieler & Hahn, 2007), DTD-Synthetic (Aota et al., 2023), AITEX (Silvestre-Blanes et al., 2019), SDD (Tabernik et al., 2020), BTAD (Mishra et al., 2021), MPDD (Jezek et al., 2021), ELPV(Deitsch et al., 2019)) and ten medical anomaly detection datasets (BrainMRI (Salehi et al., 2021), HeadCT (Salehi et al., 2021), LAG (Li et al., 2019a), Br35H (Hamada, ), CVC-ColonDB (Tajbakhsh et al., 2015), CVC-ClinicDB (Bernal et al., 2015), Kvasir (Jha et al., 2020), Endo (Hicks et al., 2021), ISIC (Gutman et al., 2016), TN3K (Gong et al., 2021)).
To assess the ZSAD performance, the full dataset of MVTec AD, including both training set and test set, is used as the auxiliary training data, on which AD models are trained, and they are subsequently evaluated on the test set of the other 18 datasets without any further training. We train the model on the full dataset of VisA when evaluating the performance on MVTec AD. Table 5 provides the data statistics of MVTec AD and VisA, while Table 6 shows the test set statistics of the other 17 datasets.
Table 5: Data statistics of MVTec AD and VisA.
Dataset Subset Type Original Training Original Test
Normal Normal Anomalous
MVTec AD Carpet Texture 280 28 89
Grid Texture 264 21 57
Leather Texture 245 32 92
Tile Texture 230 33 83
Wood Texture 247 19 60
Bottle Object 209 20 63
Capsule Object 219 23 109
Pill Object 267 26 141
Transistor Object 213 60 40
Zipper Object 240 32 119
Cable Object 224 58 92
Hazelnut Object 391 40 70
Metal_nut Object 220 22 93
Screw Object 320 41 119
Toothbrush Object 60 12 30
VisA candle Object 900 100 100
capsules Object 542 60 100
cashew Object 450 50 100
chewinggum Object 453 50 100
fryum Object 450 50 100
macaroni1 Object 900 100 100
macaroni2 Object 900 100 100
pcb1 Object 904 100 100
pcb2 Object 901 100 100
pcb3 Object 905 101 100
pcb4 Object 904 101 100
pipe_fryum Object 450 50 100
Table 6: Data statistics of the other 17 AD datasets. They are used for ZSAD inference only.
Data type Dataset Modalities |C| Normal Anomalous
Object SDD Photography 1 286 54
BTAD Photography 3 451 290
MPDD Photography 6 176 282
Textual AITEX Photography 12 564 183
DAGM Photography 10 6996 1054
DTD-Synthetic Photography 12 357 947
ELPV Electroluminescence 2 377 715
Brain BrainMRI Radiology (MRI) 1 98 155
HeadCT Radiology (CT) 1 100 100
Br35H Radiology (MRI) 1 1500 1500
Fundus LAG Fundus Photography 1 786 1711
Colon CVC-ColonDB Endoscopy 1 0 380
CVC-ClinicDB Endoscopy 1 0 612
Kvasir Endoscopy 1 0 1000
Endo Endoscopy 1 0 200
Skin ISIC Photography 1 0 379
Thyroid TN3K
Radiology (Utralsound)
1 0 614
Appendix B Implementation Details
B.1 Details of Model Configuration.
Following previous works (Deng et al., 2023; Chen et al., 2023a; Zhou et al., 2024), FAPrompt adopts a modified version of CLIP –OpenCLIP (Ilharco et al., 2021) and its publicly available pre-trained backbone VIT-L/14@336px– as the VLM backbone to enhance the model’s attention to local features while preserving its original structure. Following Zhou et al. (2024), we replace the original Q-K self-attention mechanism in the visual encoder with a V-V self-attention mechanism during patch feature extraction, starting from the 6th layer of the visual encoder. The parameters of both the visual and text encoders in CLIP are frozen throughout the experiments.
Inspired by previous works (Jia et al., 2022; Zhou et al., 2024; Khattak et al., 2022), We use text prompt tuning to refine the original textual space of CLIP by adding additional learnable token embeddings into its text encoder. By default, the learnable token embeddings are attached to the first 9 layers of the text encoder to refine the textual space, with a token length of four for each layer. The lengths of the learnable normal prompt and abnormal tokens in CAP are set to five and two, respectively. The number of fine-grained abnormality prompts ($K$) and selected patch tokens ($M$) in DAP are both set to 10. To align with the dimension of VIT-L/14@336px, the abnormality prior network $\psi(\cdot)$ is configured with the input and output dimensions of $768\times M$ and 768, respectively, and includes a hidden layer of size $(768\times M)/16$ with ReLU activation.
We utilize the Adam optimizer with an initial learning rate of 1e-3 to update the model parameters. The input images are resized to 518$\times$518 with a batch size of eight. This resizing is also applied to other baseline models for a fair comparison, while preserving their original data preprocessing methods, if applicable. The training is conducted for seven epochs across all experiments. During the inference stage, a Gaussian filter with $\sigma=10$ is applied to smooth the anomaly score map. All experiments are conducted using PyTorch on a single GPU (NVIDIA GeForce RTX 3090).
B.2 Implementation of Comparison Methods
To evaluate the efficiency of FAPrompt, we compare its performance against eight state-of-the-art (SotA) baselines. The results for CLIP (Ilharco et al., 2021), CLIP-AC (Ilharco et al., 2021), WinCLIP (Jeong et al., 2023), APRIL-GAN (Chen et al., 2023a), CoOp (Zhou et al., 2022b), and AnomalyCLIP (Zhou et al., 2024) are sourced from AnomalyCLIP, except the newly added datasets (SDD, AITEX, ELPV, LAG). For fair comparison, these implementations follow the setup of AnomalyCLIP. We use the official implementations of AnoVL (Deng et al., 2023) and CoCoOp (Zhou et al., 2022a). To adapt CoCoOp for ZSAD, we replace its learnable text prompt templates with normality and abnormality text prompt templates, which is consistent with the implementation of CoOp in existing ZSAD studies. All other parameters remain consistent with those specified in their original papers.
Appendix C Additional Results
C.1 Model Complexity of FAPrompt vs. SotA Methods
We compare the model complexity of FAPrompt with SotA methods in Table 7, evaluating the number of parameters, per-batch training time, and per-image inference time. The batch size for all approaches is set to eight for fair comparison, excluding training-free methods WinCLIP and AnoVL. While FAPrompt introduces additional trainable parameters, leading to a slightly longer training time, this minor computational overhead results in substantial performance improvements over competing methods. Additionally, since training is performed offline, this training computational overhead is generally negligible in real-world applications. In terms of inference time, our approach remains reasonably efficient and responsive.
Table 7: Number of parameters, per-batch training time (ms) and per-image inference time (ms) in comparison with competing methods.
Model Number of Para. Training Time Inference Time
WinCLIP 0 0 227.5±0.7
AnoVL 0 0 171.4±0.5
APRIL-GAN 3148800 368.7±0.5 47.9±0.1
CoOp 9216 643.8±1.1 89.9±0.7
CoCoOp 83760 737.4±3.6 93.8±0.7
AnomalyCLIP 5555200 914.1±0.9 124.2±0.9
FAPrompt 9612256 1354.1±1.7 214.7±0.8
C.2 Comparison with SOTA Full-shot Methods
We also conduct experiments on five of the most commonly used datasets to examine the performance gap between FAPrompt and two SotA full-shot methods, PatchCore (Roth et al., 2022) and RD4AD (Deng & Li, 2022). Note that it is not a fair comparison as PatchCore and RD4AD utilize the full training data of each testing dataset in its detection while ZSAD methods like FAPrompt does not use any of such training data. The results presented in Table 8 are only for analyzing the possible upper bound performance of ZSAD. Despite the unfair utilization of the dataset-specific training data in PatchCore and RD4AD, FAPrompt obtains rather impressive detection performance, further reducing the performance gap between ZSAD and full-shot methods.
Table 8: Comparison of ZSAD performance between FAPrompt and two SotA full-shot methods. The best and second-best results are respectively highlighted in red and blue.
Dataset AnomalyCLIP FAPrompt PatchCore RD4AD
Image-level (AUROC, AP)
MVTecAD (91.5, 96.2) (91.9, 95.7) ( 99.0, 99.7) (98.7, 99.4)
VisA (82.1, 85.4) (84.5, 86.8) (94.6, 95.9) ( 95.3, 95.7)
BTAD (88.3, 87.3) (92.0, 92.2) (93.2, 98.6) ( 93.8, 96.8)
MPDD (77.0, 82.0) (80.6, 83.3) ( 94.1, 96.3) (91.6, 93.8)
DAGM (97.5, 92.3) ( 98.9, 95.7) (92.7, 81.3) (92.9, 79.1)
Pixel-level (AUROC, PRO)
MVTecAD (91.1, 81.4) (90.6, 83.3) ( 98.1, 92.8) (97.8, 93.6)
VisA (95.5, 87.0) (95.9, 87.5) ( 98.5, 92.2) (98.4, 91.2)
BTAD (94.2, 74.8) (95.6, 75.2) (97.4, 74.4) ( 97.5, 75.1)
MPDD (96.5, 87.0) (96.5, 87.9) ( 98.8, 94.9) (98.4, 95.2)
DAGM (95.6, 91.0) ( 98.3, 95.4) (95.9, 87.9) (96.8, 91.9)
C.3 t-SNE Visualization of Prompt-wise Anomaly Scores
To explore the complementarity of abnormality prompts in FAPrompt, we visualize the two-dimensional t-SNE results of the anomaly score map $S_{x}^{a}$ across various categories in the VisA dataset. Fig. 6 presents such results on three more datasets from VisA to substantiate the analysis we discussed based on Fig. 3.
Figure 6: t-SNE Visualization of prompt-wise anomaly scores on diverse categories in VisA.
C.4 Hyperparameter Sensitivity Analysis
We present the pixel-level results for the sensitivity w.r.t. the number of abnormality prompts ($K$) in CAP and the number of selected patch tokens ($M$) in DAP in Fig. 7. The trend of the results is consistent with the image-level performance shown in Fig. 4.
Figure 7: Averaged pixel-level ZSAD results on industrial and medical data with varying $K$ and $M$.
C.5 Qualitative Results of FAPrompt
We compare the anomaly maps generated by FAPrompt with those produced by other ZSAD models across various datasets, as shown in Fig. 8. APRIL-GAN and AnomalyCLIP are selected as representatives of handcrafted and learnable text prompt competitors, respectively. The visualization results show that FAPrompt demonstrates significantly more accurate segmentation compared to the other two methods across both industrial and medical domains. In particular, despite not accessing any additional information or training from medical data, FAPrompt effectively localizes abnormal lesion/tumor regions, which highlight the cross-dataset generalization superiority of the fine-grained abnormality semantics learned by FAPrompt.
Figure 8: Visualization of anomaly maps generated by different ZSAD methods.
We also provide additional pixel-level anomaly score maps on diverse datasets to further showcase the strong segmentation capability of FAPrompt in Figs. 9 to 18. Specifically, for the industrial AD datasets, we select three object categories (capsule, pipe_fryum in VisA and metal_plate in MPDD) and three texture categories (grid, tile in MVTecAD and AITEX) for visualization. For the medical AD datasets, we visualize the pixel-level anomaly detection performance for the brain, colon, skin, and thyroid anomalies.
Figure 9: Anomaly maps generated by FAPrompt for the capsules category in VisA. The first row represents the input images, while the second row displays the ground truth of anomalous regions. The bottom row illustrates the segmentation results from FAPrompt.
Figure 10: Anomaly maps generated by FAPrompt for the pipe_fryum category in VisA. The first row represents the input images, while the second row displays the ground truth of anomalous regions. The bottom row illustrates the segmentation results from FAPrompt.
Figure 11: Anomaly maps generated by FAPrompt for the metal_plate category in MPDD. The first row represents the input images, while the second row displays the ground truth of anomalous regions. The bottom row illustrates the segmentation results from FAPrompt.
Figure 12: Anomaly maps generated by FAPrompt for grid category in MVTecAD. The first row represents the input images, while the second row displays the ground truth of anomalous regions. The bottom row illustrates the segmentation results from FAPrompt.
Figure 13: Anomaly maps generated by FAPrompt for tile category in MVTecAD. The first row represents the input images, while the second row displays the ground truth of anomalous regions. The bottom row illustrates the segmentation results from FAPrompt.
Figure 14: Anomaly maps generated by FAPrompt for AITEX. The first row represents the input images, while the second row displays the ground truth of anomalous regions. The bottom row illustrates the segmentation results from FAPrompt.
Figure 15: Anomaly maps generated by FAPrompt for brain-related anomalies. The first row represents the input images, while the second row displays the ground truth of anomalous regions. The bottom row illustrates the segmentation results from FAPrompt.
Figure 16: Anomaly maps generated by FAPrompt for colon-related anomalies. The first row represents the input images, while the second row displays the ground truth of anomalous regions. The bottom row illustrates the segmentation results from FAPrompt.
Figure 17: Anomaly maps generated by FAPrompt for skin-related anomalies. The first row represents the input images, while the second row displays the ground truth of anomalous regions. The bottom row illustrates the segmentation results from FAPrompt.
Figure 18: Anomaly maps generated by FAPrompt for thyroid-related anomalies. The first row represents the input images, while the second row displays the ground truth of anomalous regions. The bottom row illustrates the segmentation results from FAPrompt.
C.6 Failure Cases and Limitations
While the proposed FAPrompt demonstrates promising detection results across various categories without any dataset-specific references, it may fail in certain cases. Fig. 19 illustrates some of these failure cases. Some cases can be attributed to annotation errors. For example, images that contain multiple types of anomalies but are only partially labeled may lead to segmentation errors due to labeling inconsistencies, as can be seen in the stain defect in Fig. 19 (1). Additionally, instrument artifacts in some medical datasets are often misinterpreted as anomalies, leading to incorrect detection, e.g., Fig. 19 (2). In other cases, FAPrompt may fail in challenging cases like the ones illustrated in Fig. 19 (3)-(6), where the anomalous regions may be too small, subtle, or overshadowed by other suspicious areas (according to FAPrompt’s interpretation). Nevertheless, as demonstrated in this figure and Figs. 9 to 18, FAPrompt consistently strives to identify the most likely abnormal regions, without relying on any reference from the target datasets. Moving forward, incorporating more prior knowledge, e.g., from in-context examples, knowledge graphs, or Large Language Models (LLMs), would be helpful for providing more discriminative information for achieving more accurate anomaly detection.
Figure 19: Failure cases of FAPrompt. The first row represents the input images, while the second row displays the ground truth of anomalous regions. The bottom row illustrates the segmentation results from FAPrompt.
Appendix D Detailed Empirical Results
D.1 Breakdown Results on VisA and MVTec AD
Tables 9 to  16 present detailed downbreak ZSAD results of FAPrompt against eight SotA methods across each category of the MVTecAD and VisA datasets.
Table 9: Breakdown AUROC results of image-level ZSAD performance comparison on MVTecAD.
Handcrafted Text Prompting Learnable Text Prompting
Data Subset CLIP CLIP-AC WinCLIP APRIL-GAN AnoVL CoOp CoCoOp AnomalyCLIP FAPrompt
Carpet 11.5 10.7 95.4 98.4 - 6.7 96.7 98.8 99.0
Grid 8.7 11.9 82.2 95.8 - 7.8 89.8 97.3 96.9
Leather 9.9 5.6 96.7 99.1 - 11.7 98.5 98.6 98.5
Tile 49.9 39.1 77.6 92.7 - 41.7 87.4 94.6 95.7
Wood 45.7 42.4 93.4 95.8 - 31.4 94.5 96.5 96.4
Bottle 17.5 23.3 89.5 83.4 - 23.1 89.7 90.4 90.3
Capsule 50.9 49.1 86.9 92.0 - 35.5 80.1 95.8 95.2
Pill 55.8 60.8 80.0 76.2 - 46.5 78.7 92.0 90.5
Transistor 51.1 48.5 74.7 62.4 - 50.1 66.2 71.0 69.8
Zipper 51.5 44.7 91.6 91.1 - 33.4 92.0 91.4 91.8
Cable 37.4 37.5 77.0 72.3 - 49.7 73.3 78.9 79.5
Hazelnut 25.2 34.0 94.3 96.1 - 30.2 95.9 97.1 97.5
Metal_nut 43.9 53.6 61.0 65.4 - 49.3 71.0 74.4 71.4
Screw 80.1 76.4 89.6 97.8 - 17.0 98.3 97.5 97.4
Toothbrush 36.3 35.0 86.9 95.8 - 64.9 89.1 91.9 89.7
MEAN 38.4 38.2 85.1 87.6 89.8 33.3 86.7 91.1 90.6
Table 10: Breakdown AP results of image-level ZSAD performance comparison on MVTecAD.
Handcrafted Text Prompting Learnable Text Prompting
Data Subset CLIP CLIP-AC WinCLIP APRIL-GAN AnoVL CoOp CoCoOp AnomalyCLIP FAPrompt
Carpet 2.9 1.9 84.1 48.5 - 0.5 94.1 90.1 94.1
Grid 0.9 2.4 57.0 31.6 - 1.0 74.5 75.6 81.6
Leather 0.2 0.0 91.1 72.4 - 1.8 97.9 92.2 95.7
Tile 21.5 16.3 51.2 26.7 - 10.1 76.9 87.6 89.3
Wood 13.7 10.3 74.1 31.1 - 5.1 93.1 91.2 92.3
Bottle 1.4 4.9 76.4 45.6 - 4.5 79.4 80.9 81.0
Capsule 13.2 14.9 62.1 51.3 - 5.7 82.8 87.2 83.9
Pill 6.0 8.2 65.0 65.4 - 3.2 84.4 88.2 87.6
Transistor 15.3 11.2 43.4 21.3 - 9.3 51.5 58.1 59.0
Zipper 17.7 15.2 71.7 10.7 - 11.6 78.3 65.3 75.1
Cable 7.3 6.9 42.9 25.7 - 12.2 55.5 64.4 68.2
Hazelnut 2.8 9.4 81.6 70.3 - 4.7 89.2 92.4 93.3
Metal_nut 2.9 10.3 31.8 38.4 - 7.0 71.5 71.0 70.9
Screw 57.8 56.2 68.5 67.1 - 6.4 93.8 88.0 89.7
Toothbrush 5.8 5.2 67.7 54.5 - 16.6 71.6 88.5 87.3
MEAN 11.3 11.6 64.6 44.0 76.2 6.6 79.6 81.4 83.3
Table 11: Breakdown AUROC results of pixel-level ZSAD performance comparison on MVTecAD.
Handcrafted Text Prompting Learnable Text Prompting
Data Subset CLIP CLIP-AC WinCLIP APRIL-GAN AnoVL CoOp CoCoOp AnomalyCLIP FAPrompt
Carpet 96.0 93.1 100.0 99.5 - 99.9 98.7 100.0 100.0
Grid 72.5 63.7 98.8 86.3 - 94.7 87.7 97.0 97.9
Leather 99.4 99.5 100.0 99.7 - 99.9 98.5 99.8 99.9
Tile 88.5 89.0 100.0 99.9 - 99.7 99.4 100.0 99.7
Wood 94.0 94.9 99.4 99.0 - 97.7 44.4 96.8 98.0
Bottle 45.9 46.1 99.2 92.0 - 87.7 80.2 89.3 89.8
Capsule 71.4 68.8 72.9 79.9 - 81.1 84.2 89.9 92.4
Pill 73.6 73.8 79.1 80.5 - 78.6 83.3 81.8 89.6
Transistor 48.8 51.2 88.0 80.8 - 92.2 77.3 92.8 81.7
Zipper 60.1 36.1 91.5 89.6 - 98.8 54.5 98.5 98.4
Cable 58.1 46.6 86.5 88.4 - 56.7 29.6 69.8 74.7
Hazelnut 88.7 91.1 93.9 89.6 - 93.5 11 97.2 96.5
Metal_nut 62.8 63.4 97.1 68.4 - 85.3 81.3 93.6 89.7
Screw 78.2 66.7 83.3 84.9 - 88.9 59 81.1 85.0
Toothbrush 73.3 89.2 88.0 53.8 - 77.5 88.6 84.7 85.6
MEAN 74.1 71.5 91.8 86.2 92.5 88.8 71.8 91.5 91.9
Table 12: Breakdown PRO results of pixel-level ZSAD performance comparison on MVTecAD.
Handcrafted Text Prompting Learnable Text Prompting
Data Subset CLIP CLIP-AC WinCLIP APRIL-GAN AnoVL CoOp CoCoOp AnomalyCLIP FAPrompt
Carpet 98.8 97.8 100.0 99.8 - 100.0 99.6 100.0 100.0
Grid 87.1 83.9 99.6 94.9 - 98.1 95.8 99.1 99.3
Leather 99.8 99.8 100.0 99.9 - 100.0 99.3 99.9 100.0
Tile 95.9 96.2 100.0 100.0 - 99.9 99.8 100.0 99.9
Wood 97.9 98.3 99.8 99.7 - 99.4 68.2 99.2 99.4
Bottle 78.9 79.8 99.8 97.7 - 96.4 93.1 97.0 96.7
Capsule 92.1 90.9 91.5 95.5 - 95.7 96.5 97.9 98.4
Pill 93.4 93.6 95.7 96.0 - 94.2 96.2 95.4 97.9
Transistor 48.1 49.9 87.1 77.5 - 90.2 71.1 90.6 78.9
Zipper 87.4 73.9 97.5 97.1 - 99.7 86.7 99.6 99.5
Cable 70.8 64.3 91.2 93.1 - 69.4 50.8 81.4 82.9
Hazelnut 94.6 95.9 96.9 94.8 - 96.7 45.9 98.6 98.1
Metal_nut 87.7 89.2 99.3 91.9 - 96.3 93.6 98.5 97.5
Screw 91.4 86.6 93.1 93.6 - 96.2 81.2 92.5 93.6
Toothbrush 90.7 96.0 95.6 71.5 - 90.4 95.1 93.7 93.8
MEAN 87.6 86.4 96.5 93.5 96.7 94.8 84.9 96.2 95.7
Table 13: Breakdown AUCROC results of image-level ZSAD performance comparison on VisA.
Handcrafted Text Prompting Learnable Text Prompting
Data Subset CLIP CLIP-AC WinCLIP APRIL-GAN AnoVL CoOp CoCoOp AnomalyCLIP FAPrompt
candle 37.9 33.0 95.7 83.8 - 46.2 63.7 79.3 87.2
capsules 69.7 75.3 85.0 61.2 - 77.2 69.8 81.5 91.6
cashew 69.1 72.7 92.2 87.3 - 75.7 93.3 76.3 90.5
chewinggum 77.5 76.9 95.3 96.4 - 84.9 96.5 97.4 97.6
fryum 67.2 60.9 75.3 94.3 - 80.0 76.6 93.0 96.5
macaroni1 64.4 67.4 77.8 71.6 - 53.6 68.0 87.2 83.1
macaroni2 65.0 65.7 66.7 64.6 - 66.5 75.4 73.4 71.4
pcb1 54.9 43.9 79.8 53.4 - 24.7 81.5 85.4 68.2
pcb2 62.6 59.5 52.6 71.8 - 44.6 61.6 62.2 66.4
pcb3 52.2 49.0 70.2 66.8 - 54.4 66.4 62.7 68.6
pcb4 87.7 89.0 84.5 95.0 - 66.0 93.8 93.9 95.4
pipe_fryum 88.8 86.4 69.4 89.9 - 80.1 91.0 92.4 97.4
MEAN 66.4 65.0 78.7 78.0 79.2 62.8 78.1 82.1 84.5
Table 14: Breakdown AP results of image-level ZSAD performance comparison on VisA.
Handcrafted Text Prompting Learnable Text Prompting
Data Subset CLIP CLIP-AC WinCLIP APRIL-GAN AnoVL CoOp CoCoOp AnomalyCLIP FAPrompt
candle 42.9 40.0 96.1 86.9 - 52.9 67.7 81.1 89.7
capsules 81.0 84.3 91.0 74.3 - 85.3 81.9 88.7 96.2
cashew 83.4 86.1 96.5 94.1 - 87.1 96.8 89.4 95.9
chewinggum 90.4 90.2 97.9 98.4 - 93.1 98.6 98.9 99.1
fryum 82.0 76.6 88.1 97.2 - 90.2 89.6 96.8 98.4
macaroni1 56.8 58.7 77.7 70.9 - 52.3 73.0 86.0 82.5
macaroni2 65.0 65.8 63.3 63.2 - 62.2 72.2 72.1 68.5
pcb1 56.9 48.4 81.8 57.2 - 36.0 82.4 87.0 72.5
pcb2 63.2 59.8 50.4 73.8 - 47.3 64.6 64.3 68.2
pcb3 53.0 47.6 70.4 70.7 - 54.8 71.1 70.0 76.5
pcb4 88.0 90.6 81.5 95.1 - 66.3 94.0 94.4 95.6
pipe_fryum 94.6 93.7 82.1 94.8 - 89.7 95.1 96.3 98.6
MEAN 71.4 70.2 81.4 81.4 81.7 68.1 82.3 85.4 86.8
Table 15: Breakdown AUROC results of pixel-level ZSAD performance comparison on VisA.
Handcrafted Text Prompting Learnable Text Prompting
Data Subset CLIP CLIP-AC WinCLIP APRIL-GAN AnoVL CoOp CoCoOp AnomalyCLIP FAPrompt
candle 33.6 50.0 88.9 97.8 - 16.3 97.9 98.8 98.9
capsules 56.8 61.5 81.6 97.5 - 47.5 89.7 95.0 96.3
cashew 64.5 62.5 84.7 86.0 - 32.5 85.8 93.8 95.2
chewinggum 43.0 56.5 93.3 99.5 - 3.4 98.5 99.3 99.3
fryum 45.6 62.7 88.5 92.0 - 21.7 93.3 94.6 94.4
macaroni1 20.3 22.9 70.9 98.8 - 36.8 98.6 98.3 98.2
macaroni2 37.7 28.8 59.3 97.8 - 27.5 99.0 97.6 96.8
pcb1 57.8 51.6 61.2 92.7 - 19.8 90.4 94.1 96.0
pcb2 34.7 38.4 71.6 89.7 - 22.9 89.3 92.4 92.7
pcb3 54.6 44.6 85.3 88.4 - 18.0 91.3 88.4 88.2
pcb4 52.1 49.9 94.4 94.6 - 14.0 93.6 95.7 97.1
pipe_fryum 58.7 44.7 75.4 96.0 - 29.2 96.1 98.2 98.1
MEAN 46.6 47.8 79.6 94.2 89.9 24.1 93.6 95.5 95.9
Table 16: Breakdown PRO results of pixel-level ZSAD performance comparison on VisA.
Handcrafted Text Prompting Learnable Text Prompting
Data Subset CLIP CLIP-AC WinCLIP APRIL-GAN AnoVL CoOp CoCoOp AnomalyCLIP FAPrompt
candle 3.6 6.0 83.5 92.5 - 1.1 92.4 96.2 95.8
capsules 15.8 22.4 35.3 86.7 - 18.4 72.8 78.5 84.9
cashew 9.6 10.9 76.4 91.7 - 1.7 93.6 91.6 90.0
chewinggum 17.8 30.2 70.4 87.3 - 0.1 86.1 91.2 90.1
fryum 12.1 29.3 77.4 89.7 - 2.6 91.3 86.8 87.1
macaroni1 8.1 13.4 34.3 93.2 - 18.1 93.9 89.8 89.9
macaroni2 20.9 18.4 21.4 82.3 - 2.7 89.5 84.2 80.3
pcb1 11.7 12.5 26.3 87.5 - 0.1 82.1 81.7 87.3
pcb2 12.8 13.9 37.2 75.6 - 0.7 72.9 78.9 77.8
pcb3 31.7 23.6 56.1 77.8 - 0.0 84.6 77.1 77.8
pcb4 17.1 20.3 80.4 86.8 - 0.0 84.8 91.3 91.7
pipe_fryum 16.7 6.0 82.3 90.9 - 0.6 96.2 96.8 97.2
MEAN 14.8 17.2 56.8 86.8 71.2 3.8 86.7 87.0 87.5
D.2 Dataset-specific Results on Ablation Study
In this section, we present the dataset-specific image-level and pixel-level ZSAD results for module ablation in Table 17 and Table 18, respectively.
Table 17: Dataset-specific image-level ZSAD results (AUROC, AP) of our ablation study.
[…]
Object VisA (82.1, 85.4) (83.8, 86.7) (83.8, 86.7) (82.7, 85.0) (81.0, 83.3) (84.5, 86.8)
SDD (98.1, 93.4) (98.6, 96.1) (98.0, 95.8) (98.1, 95.5) (98.3, 95.3) (98.6, 95.9)
BTAD (88.3, 87.3) (91.5, 92.4) (90.8, 91.1) (90.7, 90.7) (91.0, 89.3) (92.0, 92.2)
MPDD (77.0, 82.0) (78.7, 81.3) (77.9, 81.3) (74.6, 78.3) (73.4, 77.8) (80.6, 83.3)
Textual AITEX (62.2, 40.4) (72.8, 55.8) (72.7, 75.4) (73.6, 54.1) (75.9, 57.8) (71.9, 53.2)
DAGM (97.5, 92.3) (97.9, 93.0) (97.9, 93.0) (96.5, 88.2) (95.7, 89.6) (98.9, 95.7)
DTD-Synthetic (93.5, 97.0) (96.3, 98.5) (95.7, 93.9) (96.0, 98.0) (96.3, 98.1) (95.9, 98.3)
ELPV (81.5, 91.3) (84.8, 92.6) (80.8, 90.7) (83.0, 91.6) (80.6, 89.9) (83.5, 92.0)
Medical BrainMRI (90.3, 92.2) (95.2, 95.2) (95.0, 94.6) (95.9, 96.0) (95.9, 96.5) (95.5, 95.6)
HeadCT (93.4, 91.6) (94.7, 94.6) (93.7, 90.4) (92.3, 90.4) (92.0, 91.0) (94.8, 93.5)
LAG (74.3, 84.9) (75.2, 85.4) (75.2, 85.4) (75.2, 85.5) (74.5, 84.6) (75.6, 85.4)
Br35H (94.6, 94.7) (97.4, 97.1) (97.1, 96.8) (97.3, 97.1) (97.0, 96.9) (97.8, 97.5)
Table 18: Dataset-specific pixel-level ZSAD results (AUROC, PRO) of our ablation study.
[…]
Object VisA (95.5, 87.0) (95.1, 85.1) (95.1, 85.0) (95.8, 86.1) (95.6, 85.1) (95.9, 87.5)
SDD (98.1, 95.2) (98.3, 93.8) (98.3, 93.2) (97.9, 95.6) (97.7, 92.5) (98.3, 93.6)
BTAD (94.2, 74.8) (94.4, 70.5) (94.4, 70.5) (95.4, 73.7) (95.5, 75.2) (95.6, 75.2)
MPDD (96.5, 87.0) (95.9, 86.2) (95.9, 86.2) (95.8, 86.4) (95.5, 85.4) (96.5, 87.9)
Textual AITEX (83.0, 66.5) (82.3, 64.5) (81.3, 61.9) (82.4, 65.2) (82.0, 62.1) (82.0, 62.6)
DAGM (95.6, 91.0) (98.1, 95.2) (97.5, 95.2) (98.5, 96.0) (98.2, 94.4) (98.3, 95.4)
DTD-Synthetic (97.9, 92.3) (97.9, 92.3) (97.9, 92.3) (98.1, 91.4) (98.1, 91.3) (98.3, 93.1)
Medical CVC-ColonDB (81.9, 71.3) (83.7, 72.8) (82.9, 68.1) (83.8, 73.9) (84.0, 73.0) (84.6, 74.7)
CVC-ClinicDB (82.9, 67.8) (83.2, 67.8) (83.4, 72.9) (83.6, 68.4) (83.3, 68.3) (84.7, 70.1)
Kvasir (78.9, 45.6) (78.8, 48.1) (78.5, 48.0) (79.3, 45.5) (79.0, 45.3) (81.2, 47.8)
Endo (84.1, 63.6) (84.3, 63.4) (84.1, 63.4) (84.7, 63.8) (84.8, 64.2) (86.4, 67.2)
ISIC (89.7, 78.4) (88.7, 78.0) (88.1, 76.8) (91.0, 80.9) (91.4, 81.3) (90.9, 81.2)
TN3K (81.5, 50.4) (84.2, 52.7) (84.5, 53.4) (84.9, 56.0) (84.2, 53.5) (84.5, 54.1)